OpenStudy (mellamokatie):

How many solutions does this system have? y = x + 5 y = -5x - 1 A. one B. none C. infinite D. two

OpenStudy (mellamokatie):

@sweetburger

OpenStudy (retireed):

Graph the two functions and see how many times they intersect, if they do.

OpenStudy (mellamokatie):

What numbers do I graph?

OpenStudy (retireed):

You graph the lines... using graph paper plot each line based on the y intercepts and the slopes of the lines

OpenStudy (retireed):

Y-intercept form for a line.... y = mx + b m is the slope and b is the y intercept the y intercept is where the point on the y axis when x is zero

OpenStudy (mellamokatie):

oh ook

OpenStudy (retireed):

Or create a table of points for each function.... 

OpenStudy (sshayer):

solve them x+5=-5x-1 x+5x=-1-5 6x=-6 x=-1 y=? how many solutions you get. ?

OpenStudy (mellamokatie):

2?

OpenStudy (mathmale):

An alternative approach would be to equate these two equations to each other: this will automatically eliminate the variable y. solve the resulting equation for x. How many such solutions are there?

OpenStudy (sshayer):

there is only one value of x and one value of y. so only one point of intersection in a plane if slopes of two lines are equal then they coincide or are parallel. first degree eq in x and y represent a st. line. if two lines in a plane are not parallel then they intersect exactly at one point.

OpenStudy (mellamokatie):

I am still confused is it 2??

sammixboo (sammixboo):

Are you still wanting an answer to this??

OpenStudy (mellamokatie):

Yes

OpenStudy (mellamokatie):

@sammixboo

OpenStudy (mathmale):

sshayer typed the following to you yesterday. Did you not see his contribution? sshayer Best Response Medals 0 there is only one value of x and one value of y. so only one point of intersection

OpenStudy (mellamokatie):

So then A?

OpenStudy (mellamokatie):

@mathmale